Multiply.
7w(2w-3)(2w+3)

Answer by SHUgrad05(58)

Multiply.
7w(2w-3)(2w+3)
7w*2w-7w*3 = 14w-21w = -7w
-7w*(2w+3)
-14w-21w = -35w